# Auto Arena of LLMs: Automating LLM Evaluations with Agent Peer-battles and Committee Discussions

### v1

@@ 4.2 Results: Alignment with Human Preferences
Table 1: Correlation analysis on evaluation benchmarks on 7 LLMs.
@@
MMLU to Arena 89.3%
OpenLLM to Arena 89.3%
MT-Bench to Arena 82.9%
MT-Bench-Hard to Arena 89.3%
LC-AlpacaEval to Arena 90.0%
Auto-Arena to Arena 96.4%
We regard Chatbot Arena (referred to as “Arena” in the following tables) scores as a trustworthy indicator of human preferences and the general capabilities of LLMs. Table 1 shows the Spearman correlations with Chatbot Arena scores achieved by various benchmarks. As all benchmarks evaluate only in English, we compare with the English-only Chatbot Arena scores. As a result, we observe that both static and model-based baselines result in a similar level of Spearman correlation around 90%, with LC-AlpacaEval slightly surpassing other benchmarks. Then, Auto-Arena is able to improve the correlation to 96.4%, outperforming SOTA performance by 6.4%. Notably, among all benchmarks, Auto-Arena is the only one that requires no human efforts, eliminating the need for manual dataset collection. The high alignment with human preferences could originate from the human-like design, which effectively mimics the human users’ voting processes.
4.3 Analysis on Peer Battles
@@
Without Peer Battles With Peer Battles
Auto-Arena to Arena 50.0% 96.4%
To investigate the effect of "peer battles" on the overall evaluation quality, we conduct an ablation study. As a baseline, we simply ask the committee to collectively evaluate the two candidates’ initial responses to the raised question. The procedure would be similar to model-based metrics such as MT-Bench and AlpacaEval. However, as we only included 40 questions, the evaluation is expected to be less reliable. The ablation results are shown in Table 2. Without the peer battle component, the Spearman correlation with human preferences drops from 96.4% to 50.0%. This result proves the importance of the peer battles, during which the performance gaps between candidates become more visible and robust to judges. Thus, peer battles can improve evaluation reliability and alignment with human preferences.
4.4 Analysis on Committee Discussions
Figure 3: Cohen’s Kappa Agreement Among Judges Before Committee Discussions.
Figure 4: Cohen’s Kappa Agreement Among Judges After One Round of Committee Discussion.
The committee discussion component is designed to introduce diverse viewpoints and produce more consistent verdicts. As shown in Figure 4, the Cohen’s Kappa agreement [21] among judges before committee discussion is very low, averaging 0.10, which indicates slight agreement. We notice that weak model judges and strong model judges has an especially low agreement, such as GPT-4 and Llama -2. This shows that general model capabilities could result in significant performance gaps when used as judges. After one round of Committee Discussion, Figure 4 shows an overall improvement in agreement across all judge combinations. In the discussion process, judges are exposed to more viewpoints, among which some may be convincing enough to result in a change in verdict. After discussions, the average Cohen’s Kappa increases to 0.38, indicating fair agreement.
@@ 4.4 Analysis on Committee Discussions
Table 3: Probability of Agreement among Judges. Agreement is defined as the mean probability of two random judges agreeing with each other.
Auto Arena Auto Arena MT-Bench
(Before discussion) (After discussion) Human evaluation
Agreement 48% 68% 67%
@@
Table 3 shows the agreement probability among judges. Agreement probability is defined as the mean probability of two random judges agreeing with each other. After committee discussion, the agreement increases by 20%, matching the agreement level among human annotators on MT-Bench. This observation indicates that committee discussions can improve the quality of judgments significantly in terms of consistency and reliability, matching human-level performance.
4.5 Adding Models to the Leaderboard
Figure 5: Elo Score Changes of Adding Llama-3 to the Existing Ranking of 7 Models.
@@ 4.5 Adding Models to the Leaderboard
Figure 7: Elo Scores of 17 Models by Auto-Arena on English.
To add a new candidate to the finished tournament, we ask it to debate with $log_{2}(n)$ opponents with similar ELO scores, where $n$ is the number of total participants after adding the new candidate. We initialize the first pairing by asking the new candidate to debate with the opponent with the most similar MMLU score.
Figure 5 shows the ELO score changes of adding a new participant (Llama-3, which was newly released) to the existing ranking. It has $log_{2}(8)=3$ models peer battles with close opponents. Firstly, it battles with Qwen-1.5 based on MMLU similarity and wins, which results in a very high ELO score, even above GPT-4. Then, it is paired with GPT-4, the most similar opponent. After losing, the score stabilizes at second place. Finally, it is paired with Claude-3-Haiku, the other most similar opponent besides GPT-4. It wins and secures the final ranking in second place. This process lets the new candidate battle with a reasonable fraction of close opponents and makes the final ranking stable without disrupting the other participants. Moreover, this mechanism of adding participants is generalizable and maintains the computational costs below $nlog_{2}(n)$.
Therefore, we repeat the process and add a total of 10 models to the existing tournament, resulting in a comprehensive leaderboard. We choose the best-performing and newest models from each major model family in the top 20 list on Chatbot Arena. Moreover, we add 4 mainstream Chinese LLMs that are not on Chatbot Arena, which are GLM, SenseChat, Minimax, and Wenxin. Figure 7 shows the overall ELO scores produced by Auto-Arena on the 17 models. On models included in Chatbot Arena, we could recover rankings with 92.3% accuracy.
Table 4: Correlation analysis on evaluation benchmarks on 17 LLMs after extension.
Spearman Correlation
MMLU to Arena 90.1%
OpenLLM to Arena 85.7%
MT-Bench to Arena 89.3%
MT-Bench-Hard to Arena 86.7%
LC-AlpacaEval to Arena 90.3%
Auto-Arena to Arena 94.5%
Table 4 shows the Spearman correlations on the set of 17 models after expansion. Auto-Arena remains the model most aligned with human preferences by a margin of 4%. Therefore, Auto-Arena of LLMs is generalizable and robust for maintaining a leaderboard on a large number of LLMs.
@@ 4.6 Easy Extension to Other Domains and Languages
Figure 7 shows the ELO scores derived by Auto-Arena for Chinese evaluation. We notice that the ranking differs significantly from English results. For example, SenseChat-5, an LLM specifically trained for Chinese usage, is able to improve from 10th place (in English) to 1st place (in Chinese). On the contrary, Llama-3-70b, which only uses 5% multilingual data during pretraining, drops from 4th place (in English) to 10th place (in Chinese). As Chinese evaluation benchmarks are limited, we compare to the Chinese-only leaderboard on Chatbot Arena, which constitute for 10.36% of all collected dialogues. On the 7 models that are also included in Chatbot Arena, Auto-Arena recovers their ELO scores with 92.86% correlation and restores rankings with 90.5% accuracy.
@@ 5 Investigation of LLM’s Behaviors in Competitive Peer Battles
Figure 8: LLM agents display competitive behaviors in peer battles.
Figure 9: LLM agents learn from each other in peer battles.
5.1 LLM can display competitive behaviors
The example in Figure 9 shows excerpts of a peer battle around question: “how many unique ways to arrange letters in ‘Letter’.” Candidate A (powered by Yi-34B-Chat) gives a wrong answer as it miscounts occurrences for repeated letters and miscalculates factorials. The opponent B (powered by Claude-3-Haiku) quickly and precisely points out these two issues and skillfully raised a follow-up that targets A’s weaknesses: “how about the word ‘BANANA’?” Then, A makes the same mistake as before. Seeing these results, the judge then determines that A is the better assistant. In peer battles, we see that LLM candidates efficiently understand the rules of the competitive environment and can design specific strategies to attack the opponent in order to win.
5.2 LLM candidates can improve by learning from its opponents
Figure 9 shows a roleplay example between Claude-3-Haiku (Candidate A) and Command-r-Plus (Candidate B). In the first round, A answers the question plainly while B, in addition to answering the question, also employs the appropriate speech style, which better matches the “roleplay” instructions. Then, in the rounds after, without any explicit instructions, A learns from its opponent and also incorporates the speech style. This case shows an interesting observation that, even in competitive environments, LLM candidates can display learning behaviors and improve from the interactions. Expanding upon this observation, using the interplay between LLM agents to improve performances could be a promising future paradigm of learning.
5.3 Peer-battles make the performance gaps become visible
@@
In the example shown in Figure 10, given a math question on infinite series, both candidate A (powered by Claude-3-Haiku) and candidate B (powered by GPT-4-Turbo) provide correct answers in the first round. However, as they raise follow-up questions to each other, the performance gap becomes more visible: Candidate B is able to provide a more elaborate and helpful response. Therefore, although the judges initially decided that it was a tie result, after seeing the subsequent debate rounds, they change to favoring assistant B.
This example shows that the debate process indeed pushes the candidate LLM’s capabilities to the limit, testing deeper understandings and reasoning abilities. Also shown in the previous Table 2, the peer-battle design is indispensable for a robust and comprehensive evaluation.
@@ 6 Conclusions
In this paper, we innovatively design a completely automatic evaluation framework: Auto-Arena of LLMs. By using LLM agents to generate questions, employing LLM candidates in peer battles, and evaluating responses using LLM committee discussions, Auto-Arena produces less-contaminated, robust, and trustworthy evaluation results. In the extensive experiments, Auto-Arena achieves the highest correlation with human preferences, despite requiring zero human efforts. It is easily adaptable to other domains and resources, promoting fairness of AI system evaluations. The peer battles also demonstrate several interesting LLM behaviors in competitive environments, including attacking and learning from the opponents. Along with the paper, we release a website, demos, and an actively-maintained leaderboard. We hope this work could serve as a valuable tool for LLM evaluation.
@@ A.2 Example Questions Generated
We list some of the example questions generated here:

### v2

@@ 4.2 Results: Alignment with Human Preferences
Table 1: Correlation analysis with Chatbot Arena of evaluation benchmarks on 7 LLMs.
@@
MMLU 89.3%
OpenLLM 89.3%
MT-Bench 82.9%
MT-Bench-Hard 89.3%
LC-AlpacaEval 90.0%
Auto-Arena 96.4%
We regard Chatbot Arena scores as a trustworthy indicator of human preferences and general capabilities of LLMs. Table 1 shows the Spearman correlations with Arena scores achieved by various benchmarks. As all benchmarks are evaluated only in English, we use English-only Arena scores. We see that both static and model-based baselines result in a similar level of correlation no higher than 90%, with LC-AlpacaEval slightly surpassing others. Then, Auto-Arena can improve the correlation to 96.4%, outperforming SOTA by 6.4%. Notably, among all benchmarks, Auto-Arena is the only one that requires no human efforts, either on dataset compilation or judgment generation. The high alignment with human preferences could originate from the human-like design, which effectively mimics the human users’ voting processes.
@@
Auto-Arena to Arena
Without Peer Battles 50.0%
With Peer Battles 96.4%
@@
Table 3: Agreement Probability among Judges. Agreement is defined as the mean probability of two random judges agreeing with each other.
Agreement
Auto Arena (Before discussion) 48%
Auto Arena (After discussion) 68%
MT-Bench Human evaluation 67%
4.3 Ablation Studies on Peer Battles and Committee Discussions
To investigate the effect of "peer battles" on the evaluation quality, we conduct an ablation study and show the results in Table 3. As a baseline, we ask the committee to collectively evaluate the two candidates’ initial responses to the raised question. Without the peer battle component, the correlation with human preferences drops from 96.4% to 50.0%. This proves the importance of the peer battles, during which the performance gaps between candidates become more visible and robust to judges. Thus, peer battles can improve evaluation reliability and alignment with human preferences.
Figure 3: Cohen’s Kappa Agreement with Majority Vote Before Committee Discussions.
Figure 4: Cohen’s Kappa Agreement with Majority Vote After Committee Discussions.
The committee discussion component is designed to introduce diverse viewpoints and produce more consistent verdicts. As shown in Figure 4, the Cohen’s Kappa agreement [29] between individual judges and the final result (voted) is low, averaging 0.36, which indicates fair agreement. Specifically, compared to strong models, the judgments of weak models align less with the voted result, such as Llama-2 compared to GPT-4. This shows that general model capabilities could result in significant performance gaps when used as judges. After the Committee Discussion, Figure 4 shows an overall improvement in agreement, averaging 0.56, which indicates moderate agreement. In the discussion process, judges are exposed to more viewpoints, among which some may be convincing enough to result in a change in verdict. More analysis on inter-judge agreement is provided in Appendix A.3.
@@ 4.3 Ablation Studies on Peer Battles and Committee Discussions
Table 3 shows the agreement probability among judges. Agreement probability is defined as the mean probability of two random judges agreeing with each other. After committee discussion, the agreement increases by 20%, matching the agreement level among human annotators on MT-Bench. This observation indicates that committee discussions can significantly improve the quality of judgments in terms of consistency and reliability, matching human-level performance.
5 Constructing a Leaderboard with Auto-Arena
Figure 5: Elo Score Changes of Adding Llama-3 to the Ranking of 7 Models.
Figure 6: Elo Scores of 17 Models by Auto-Arena on English.
5.1 Update New Models to Leaderboard
With Auto-Arena, we can obtain the rank for a list of models with their ELO scores to construct a leaderboard. As new LLMs come out frequently, we describe how to add new candidate models to the existing leaderboard with 10 more models that are released very recently. To add a new candidate, we ask it to debate with $\lceil log_{2}(n)\rceil$ opponents with similar ELO scores, where $n$ is the number of total participants after adding the new candidate. We initialize the first pairing by asking the new candidate to debate with the opponent with the most similar MMLU score. This addition mechanism is generalizable and maintains the computational costs of evaluating $n$ models below $nlog_{2}(n)$.
As an example, we add a new participant (Llama-3) to the existing 7-model ranking. It battles with $\lceil log_{2}(8)\rceil=3$ close opponents and Figure 6 shows the ELO score changes. Firstly, it is paired with Qwen-1.5 based on MMLU similarity and wins, which results in a very high ELO score, even above GPT-4. Then, it is paired with GPT-4, the most similar opponent. After losing, the score stabilizes at second place. Finally, it is paired with Claude-3-Haiku, the closest opponent besides GPT-4. It wins and secures the final ranking in second place. This process lets the new candidate battle with a reasonable fraction of close opponents and makes the final ranking stable without disrupting the other participants, whose score distribution remains similar before and after the addition.
Table 4: Correlation analysis with Chatbot Arena of evaluation benchmarks on 17 LLMs after extension. The 4 models not on Chatbot Arena are skipped.
Spearman Correlation
MMLU 90.1%
OpenLLM 85.7%
MT-Bench 89.3%
MT-Bench-Hard 86.7%
LC-AlpacaEval 90.3%
Auto-Arena 94.5%
@@ 5.1 Update New Models to Leaderboard
Using this scalable addition approach, we construct a comprehensive leaderboard by adding 10 models to the existing tournament. Besides the best-performing and newest models from each major model family in the top 20 list on Chatbot Arena (detailed in Appendix A.4), we add 4 popular Chinese LLMs, which are GLM [^5] https://open.bigmodel.cn/, SenseChat [^6] https://platform.sensenova.cn/home, Minimax [^7] https://platform.minimaxi.com/examination-center/text-experience-center, and Wenxin [^8] https://cloud.baidu.com/wenxin.html. Figure 6 shows the overall ELO scores by Auto-Arena on the 17 models. On the 13 models included in Chatbot Arena, we can recover rankings with 92.3% accuracy. Table 7 shows the Spearman correlations after expansion. Auto-Arena remains the method most aligned with human preferences by a margin of 4%. Therefore, Auto-Arena is generalizable and robust for maintaining a leaderboard on a large number of LLMs.
@@ 5.2 Easy Extension to Other Domains and Languages
Figure 7 shows the ELO scores derived by Auto-Arena for Chinese evaluation. We notice that the ranking differs significantly from English results. For example, SenseChat-5, an LLM specifically trained for Chinese usage, is able to improve from 10th place (in English) to 2nd place (in Chinese). On the contrary, Llama-3-70b, which only uses 5% multilingual data during pretraining [30], drops from 4th place (in English) to 10th place (in Chinese). As Chinese evaluation benchmarks are limited, we compare to the Chinese-only leaderboard on Chatbot Arena, which constitute for 10.36% of all collected dialogues. On the 7 models that are also included in Chatbot Arena, Auto-Arena recovers their ELO scores with 92.86% correlation and restores rankings with 90.5% accuracy.
Figure 8: LLM agents display competitive behaviors in peer battles.
Figure 9: LLM agents learn from each other in peer battles.
@@ 6 Investigation of LLM’s Behaviors in Competitive Peer Battles
LLMs Can Skillfully Attack the Opponents
The example in Figure 9 shows excerpts of a peer battle around question: “how many unique ways to arrange letters in ‘LETTER’.” Candidate A (powered by Yi-34B-Chat) gives a wrong answer as it miscounts occurrences for repeated letters and miscalculates factorials. The opponent B (powered by Claude-3-Haiku) quickly and precisely points out these two issues and skillfully raised a follow-up that targets A’s weaknesses: “how about the word ‘BANANA’?” Then, A still miscalculates factorials. Seeing these results, the judge then determines that B is the better assistant. In peer battles, we see that LLM candidates efficiently understand the rules of the competitive environment and can design targeted strategies to attack the opponent in order to win. The peer battle process then ensures that each raised question is tailored to the specific LLM, resulting in more specialized evaluations.
LLM Candidates Can Improve by Learning From Its Opponents
Figure 9 shows a roleplay example between Claude-3-Haiku (A) and Command R$+$ (B). In the first round, A answers the question plainly while B, in addition to answering the question, also employs the appropriate speech style, which better matches the “roleplay” instructions. Then, in the rounds after, without any explicit instructions, A learns from its opponent and also incorporates the speech style. This case shows an interesting observation that, even in competitive environments, LLM candidates can display learning behaviors and improve from the interactions. Expanding upon this observation, using the interplay between LLM agents to improve performances could be a promising future paradigm of learning.
Peer Battles Make the Performance Gaps Become Visible
@@
In the example shown in Figure 10, given a math question on infinite series, both candidate A (Claude-3-Haiku) and candidate B (GPT-4-Turbo) provide correct answers in the first round. However, as the debate deepens, the performance gap becomes more visible: Candidate B is able to provide a more elaborate and helpful response. In the ablation study without peer-battles, the judges initially decided that it was a tie. However, after seeing the subsequent debates, they change to favoring assistant B. This example shows that the debate process indeed pushes the candidate LLM’s capabilities to the limit, testing deeper understandings and reasoning abilities. Also shown in the previous Table 3, the peer-battle design is indispensable for a robust and comprehensive evaluation.
@@ 7 Conclusions
In this paper, we innovatively design a completely automatic evaluation framework: Auto-Arena of LLMs. By using LLM agents to generate questions, employing LLM candidates in peer battles, and evaluating responses using LLM committee discussions, Auto-Arena produces less-contaminated, timely, and trustworthy evaluations. In the extensive experiments, Auto-Arena achieves the highest correlation with human preferences, despite requiring zero human efforts. It is easily adaptable to other domains and resources, promoting fairness of AI system evaluations. The peer battles also demonstrate several interesting LLM behaviors in competitive environments, including attacking and learning from the opponents. Moreover, there still exist some limitations to the current approach. For example, the distribution of question domains is artificially designed, which may deviate from real-life distributions. Currently, Auto-Arena focuses on 1-to-1 peer battles, which limits its usage in multi-player scenarios. Along with the paper, we release a website, demos, and an actively-maintained leaderboard. We hope this work could serve as a valuable tool for LLM evaluation.
@@ A.2 Example Questions Generated
To show the overall quality of the questions generated, we list 2 generated questions per category here. The questions shown are not manually-selected, but simply the first 2 questions generated. The quality is consistent throughout. We manually examine the questions with closed-form answers (math, reasoning, coding) and find that all questions used are solvable.
@@ 2. What is the integral of ($x^{2}+2x+2)/(x^{3}+3x^{2}+3x+1)dx$?
A.3 Inter-judge Agreement
Figure 11: Cohen’s Kappa Agreement with Majority Vote Before Committee Discussions.
Figure 12: Cohen’s Kappa Agreement with Majority Vote After 1 Round of Committee Discussion.
As shown in Figure 4, the Cohen’s Kappa agreement [29] among judges before committee discussion is very low, averaging 0.10, which indicates slight agreement. We notice that weak model judges and strong model judges has an especially low agreement, such as GPT-4 and Llama -2. This shows that general model capabilities could result in significant performance gaps when used as judges.
A.4 Model selection for the main experiment
Table 7: Model Selection for the Main Experiment. “Newest” and “Strongest” refer to the state at the time of experiments (2024 April). Bolded models are selected for the primary experiment with 7 models. Unbolded models are the ones added during extension.
Model Name Reasons for Inclusion License
GPT-4-0409-Turbo [34] Newest and Strongest in GPT-4 model family Proprietary
Claude-3-Haiku [2] Newest and Cheapest in Claude-3 model family Proprietary
Qwen1.5-72B-chat [3] Strongest in Qwen model family Qianwen LICENSE
Mixtral-8x7b-Instruct-v0.1 [23] Strongest in open-source Mistral small models Apache 2.0
MOE Structure
Yi-34B-Chat [1] Strongest in Yi Model Family Yi License
GPT-3.5-Turbo-0125 [32] Newest in GPT-3.5 Model Family Proprietary
Llama-2-70b-chat [38] Strongest in Llama Model Family Llama 2 Community
GPT-4o-2024-05-13 [33] Newly released model in GPT-4 Model Family Proprietary
Llama-3-70b-chat-hf [30] Newly released model in Llama Model Family Llama 3 Community
Reka-core-20240501 [37] Strongest model in Reka Model Family Proprietary
Command R Plus [15] Strongest model in Command R Model Family CC-BY-NC-4.0
mistral-large-2402 [16] Strongest model in mistral Family Proprietary
deepseek-llm-67b-chat [17] Strongest model in deepseek Family DeepSeek License
Wenxin-5 Famous Chinese model, not on Chatbot Arena Proprietary
GLM-4 Famous Chinese model, not on Chatbot Arena Proprietary
Minimax-abab6.5 Famous Chinese model, not on Chatbot Arena Proprietary
SenseChat-5 Famous Chinese model, not on Chatbot Arena Proprietary
In Table 7, we show all the models selected for the main experiment and expansion. We also include the reasons for selection. Overall, we try to select a representative set of famous models on Chatbot Arena top 20 list. While the Chatbot Arena ranking mostly consists of models with different versions, we only select the strongest or newest model from each model family. Besides the models on Chatbot Arena, we include 4 under-evaluated famous Chinese models to investigate their performances.
A.5 Comparison of baseline methods and Auto-Arena
Table 8: Comparison between Auto-Arena and Other Benchmarks.
Method Manual Construction Freshness Eval. Cost Judge
of Queries per Model
MMLU [22] Yes Static - Answer Accuracy
OpenLLM Leaderboard [6] Yes Static - Answer Accuracy
LC-AlpacaEval [19] Yes Static \$10 Single LLM (GPT-4)
MT-Bench [44] Yes Static \$10 Single LLM (GPT-4)
MT-Bench Hard [26] Yes Frequent Updates \$25 Single LLM (GPT-4)
Chatbot Arena [44] Yes Live Very High Humans
Auto-Arena No Freshly Generated \$5 Committee of LLMs
Table 8 shows a comparison between benchmark evaluation methods and Auto-Arena. Compared to previous methods, the main advantage of Auto-Arena is the zero need for human dataset construction or intervention and the freshness of queries. Another main innovation compared to previous model-based systematic benchmarking procedures is using a committee of LLMs to discuss and vote for a final winner, which introduces diverse viewpoints.
For the evaluation cost, the costs of Auto-Arena are on the same scale as other benchmarks: We note that the primary experiment among 7 models costs around \$35 USD. Therefore, the estimated cost is \$5 per model. As models on the ranking board increase, the costs of conducting debates should grow slowly in log scale, which comes from conducting $nlog_{2}(n)$ pairings when adding 1 model to a ranking of $(n-1)$ models. The evaluation costs, however, shall remain the same as we use a committee of 5 LLMs at all times.